OpenStudy (anonymous):

Hey if this is in row echolen form then how about will you solve it? 1 0 8 -5 6 0 1 4 -9 3 0 0 1 1 2

OpenStudy (anonymous):

Solve what? Is it an augmented matrix for a system of linear equations?

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

What was the original system? Or was it a question where its already in this form?

OpenStudy (anonymous):

yes i've solved it..but the answer in the answer book states other wise. hold on quick

OpenStudy (anonymous):

Please post the original system

OpenStudy (anonymous):

there's no original system. the question states if this has been reduced to row echolen form then find the solution. so i know i have to make it into reduced row echolen form

OpenStudy (anonymous):

i got this as my answer 1 0 0 -13 -10 0 1 0 -13 -5 0 0 1 1 2

OpenStudy (anonymous):

and book states x1 = 13t - 10 and x2 = 13t - 5 and x3 = t - 2 and x4 = t

OpenStudy (anonymous):

was wondering how they end up getting that?

OpenStudy (anonymous):

The idea of getting it into reduced form is so that it is simple to solve the system. Add the variables back in and solve, which is now very simple.

OpenStudy (anonymous):

x_1 - 13x_4 = -10 x_2 - 13x_4 = -5 x_3 + 1x_4 = 2 x_4 is arbitrary

OpenStudy (anonymous):

Since x_4 is arbitrary we can call it t our parameter: Then we have x_1 = -10 + 13t x_2 = -5 + 13t x_3 = 2 - t

OpenStudy (anonymous):

This is the same as in your book